1. Determine the domain of f in f(x) = 10 - square root of 3 -x
2. What is the function which would intersect the y-axis 2 positions up (vertical) from the function F(x)=x^2?
3. For f(x) = 4x+1 and g(x)=x^2-2x-4 find (g-f)(x)

Answer by user_dude2008(1862)

1. 3-x >= 0 ---> x <= 3 Domain: (-infinity, 3]
2. Function is G(x)=x^2+2
3. (g-f)(x)=(x^2-2x-4)-(4x+1)=x^2-2x-4-4x-1=x^2-6x-5